# Forecasts for Galaxy Formation and Dark Matter Constraints from Dwarf Galaxy Surveys

### v1

@@ 4 Galaxy Formation Forecasts
In both cases, we set $\sigma_{M}=0.2~{}\mathrm{dex}$, and $\mathcal{S}_{\mathrm{gal}}=0.2$. The “weak” cutoff scenario is broadly consistent with predictions from high-resolution simulations of galaxy formation at high redshifts (e.g., see Côté et al. 2018) and from various SAMs (e.g., Kravtsov & Manwadkar 2022; Ahvazi et al. 2023). Meanwhile, the “strong” cutoff scenario is more consistent with predictions from hydrodynamic simulations run to lower redshifts (e.g., Sawala et al. 2016; Fitts et al. 2017; Munshi et al. 2021) and other SAMs (e.g., Benitez-Llambay & Frenk 2020). We note that galaxy occupation in hydrodynamic simulations is resolution and definition dependent (e.g., see Nadler et al. 2020b; Ahvazi et al. 2023), and that higher-resolution simulations (e.g., Wheeler et al. 2019) and convergence studies (e.g., Munshi et al. 2021) indicate that progressively lower-mass halos host “galaxies” (i.e., simulated systems with more than a critical number of star particles) as resolution increases.
@@ 4.1 Cutoff Mass Scale Posteriors
Figure 3 illustrates our results for the “strong” cutoff scenario: the left panel shows the marginalized posterior for $\mathcal{M}_{50}$ given observations of all satellites around one (blue) and two (red) MW-mass hosts, and the right panel shows the corresponding galaxy occupation fraction posteriors. We find $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})=7.9^{+0.4}_{-1.6}$ at $68\%$ confidence for one host. Thus, $\mathcal{M}_{50}$ can be probed at the $\approx 1\sigma$ level using a single complete satellite population; however, a long tail toward low $\mathcal{M}_{50}$ prevents a measurement at the $2\sigma$ level. Note that our inferred distribution of $\mathcal{M}_{50}$ is consistent with the input value of $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})=8.0$, indicating that our galaxy–halo connection reconstruction is unbiased.
The tail toward low $\mathcal{M}_{50}$ arises because the “strong” cutoff primarily suppresses satellite abundances at the lowest luminosities we consider, and this effect is not significant enough to be distinguished from models that do not suppress the abundance of satellites with $M_{V}<0~{}\mathrm{mag}$ and $\mu_{V}<32~{}\mathrm{mag\ arcsec}^{-2}$. As a result, models with $\mathcal{M}_{50}\lesssim 10^{7}~{}M_{\mathrm{\odot}}$ are not distinguishable from each other and have roughly equal probability in the marginalized posterior. Our analysis of the full one-host posterior in Appendix C.1 shows that the low-$\mathcal{M}_{50}$ tail is enhanced by a degeneracy with the luminosity scatter $\sigma_{M}$. For large $\sigma_{M}$, galaxies hosted by abundant, low-mass halos preferentially up-scatter in luminosity, resulting in higher predicted satellite abundances at all luminosities. In this regime, galaxy–halo connection parameters adjust so that lower-mass halos explain more of the mock data; in turn, large values of $\mathcal{M}_{50}$ are disfavored because they prevent these low-mass halos from forming galaxies. In other words, higher luminosity scatter increases the sensitivity of a fixed-depth survey to lower-mass subhalos, boosting the constraining power on large values of $\mathcal{M}_{50}$.
@@
Figure 3: Left: marginalized posteriors for the galaxy formation cutoff mass scale $\mathcal{M}_{50}$ (i.e., the peak virial mass at which $50\%$ of halos host galaxies with $M_{V}<0~{}\mathrm{mag}$), from our forecasted measurements of one (blue) and two (red) complete satellite populations of MW-mass hosts. Dark (light) shaded regions show $68\%$ ($95\%$) confidence intervals. Right: marginalized posteriors for the galaxy occupation fraction, defined as the fraction of halos that host galaxies with $M_{V}<0~{}\mathrm{mag}$ as a function of peak virial mass. Dark (light) bands show the $68\%$ ($95\%$) confidence intervals from our one and two-host forecasts. In both panels, black dashed lines show the input “strong” galaxy formation cutoff.
@@ 4.3 Projections for Additional Hosts
As discussed in Section 3.2, it is computationally challenging to run our full MCMC forecasts for more than two hosts. Instead, we derive projected constraints on $\mathcal{M}_{50}$ by calculating the likelihood as a function of $\mathcal{M}_{50}$ with all remaining galaxy–halo connection parameters fixed to their input values. We perform this test as a function of $N_{\mathrm{hosts}}$, sampling realizations of our galaxy–halo connection model and different host combinations for each cutoff scenario and choice of $N_{\mathrm{hosts}}$. This procedure optimistically neglects galaxy–halo connection degeneracies to illustrate the ideal scaling of $\mathcal{M}_{50}$ constraints for statistically-limited measurements; the resulting projections can be interpreted as cosmic variance-limited constraints from satellite populations.
@@
Figure 5 shows projected $1\sigma$ uncertainties on $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ as a function of $N_{\mathrm{hosts}}$ in the “strong” cutoff scenario. We derive these uncertainties from the Fisher information evaluated at the input value of $\mathcal{M}_{50}$. For large $N_{\mathrm{hosts}}$, we find that the $1\sigma$ uncertainty on $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ scales as $N_{\mathrm{hosts}}^{-1/2}$, on average, as expected when adding independent measurements. With $32$ complete hosts, we project $\approx 2\%$ uncertainties on $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ at the $1\sigma$ level, assuming negligible galaxy–halo connection systematics; different host combinations affect these projected constraints at the sub-percent level. For reference, there are roughly $30$ MW-mass hosts in the Local Volume with partially-characterized satellite populations (e.g., Carlsten et al. 2022). Our projected uncertainties do not significantly change in the “weak” cutoff scenario, although they should only be regarded as upper limits at the $1\sigma$ level in this case.
@@
In Appendix B, we show that there is a prominent tail toward low $\mathcal{M}_{50}$ in our idealized likelihoods for the “strong” and “weak” cutoff scenarios; this tail persists for large $N_{\mathrm{hosts}}$. Thus, $\mathcal{M}_{50}$ remains difficult to detect at the $2\sigma$ level even for statistically-limited measurements. Interestingly, we find that constraining power qualitatively improves if satellites that do not pass our fiducial detection thresholds are included in the inference. For example, using $M_{V}<+2~{}\mathrm{mag}$ (comparable to the luminosity of the recently-discovered Ursa Major III/UNIONS 1 dwarf galaxy candidate; Smith et al. 2023) and $\mu_{V}<34~{}\mathrm{mag\ arcsec}^{-2}$ (fainter than any known system and comparable to “stealth galaxies” hypothesized in previous studies; Bullock et al. 2010) disfavors models with low $\mathcal{M}_{50}$ due to non-detections of extremely faint systems. Observations of satellites below our fiducial detection thresholds can therefore improve our projected galaxy formation cutoff constraints; we leave a study of these systems in the context of our modeling framework to future work.
Figure 4: Marginalized posterior for $\mathcal{M}_{50}$ versus $\sigma_{M}$ from our forecast using one (blue) and two (red) complete satellite populations of MW-mass hosts in the presence of a “strong” galaxy formation cutoff. Dashed lines mark the input values of $\mathcal{M}_{50}$ and $\sigma_{M}$; two-dimensional contours represent $68$ and $95\%$ confidence intervals. Top and side panels show marginal posteriors, where shaded regions represent $68\%$ confidence intervals. The two hosts shown by the red contour have significantly different bright satellite abundances, which markedly improves the recovery of $\sigma_{M}$ and indirectly improves $\mathcal{M}_{50}$ constraints.
@@
Figure 5: Projected $1\sigma$ uncertainties for statistically-limited measurements of $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ in the “strong” cutoff scenario, holding all remaining galaxy–halo connection parameters fixed. Points show mean measurement uncertainties as a function of the number of hosts combined in the inference for $N_{\mathrm{hosts}}=1$ (blue), $2$ (red), $4$ (orange), $8$ (green), $16$ (cyan), and $32$ (gray); the gray band shows the $68\%$ range of these uncertainties due to different host combinations. The black dashed line shows the $N_{\mathrm{hosts}}^{-1/2}$ scaling expected for statistically-limited measurements when adding independent hosts.
@@ 4.4 Interpretation
Our forecasts indicate that complete measurements of two satellite populations (e.g., all satellites of the MW and M31) can probe a galaxy formation cutoff with $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$, at the $1\sigma$ level. Moreover, for statistically-limited measurements, the uncertainty of this projected $1\sigma$ constraint shrinks as additional hosts are added. We emphasize that our forecasts only rely on halos that actually host galaxies. In particular, non-detections of faint, low surface brightness galaxies—which would be hosted by halos with $M_{\mathrm{peak}}\lesssim\mathcal{M}_{50}$ in the absence of a galaxy formation cutoff—play a crucial role in our inference. In practice, carefully modeling observational incompleteness (e.g., following the efforts of Drlica-Wagner et al. 2020 and Doliva-Dolinsky et al. 2022 for the MW and M31 systems, respectively), which can mimic such non-detections, will be necessary to robustly extract a galaxy formation cutoff signal from future dwarf galaxy population data. Upcoming facilities like Rubin will make substantial progress toward these observations (Mutlu-Pakdil et al., 2021); however, we do not make specific projections for Rubin because careful modeling of, e.g., star–galaxy separation is needed to accurately characterize its dwarf galaxy detection sensitivity (K. Tsiane et al., in preparation).
Note that mock satellite populations in our analyses range from $\approx 100$ to $200$ objects per host. Given that subhalo and satellite statistics are approximately Poissonian (e.g., Mao et al. 2015), the statistical uncertainty of our measurements is $\approx 10\%$ per host. Because uncertainties on galaxy–halo connection parameters are much larger than $10\%$ in general, our projected measurements are limited by galaxy–halo connection systematics rather than statistical uncertainties. Thus, using auxiliary data to constrain galaxy–halo connection parameters at this level can reduce uncertainties in our full MCMC forecasts and help realize the statistically-limited measurements derived in Section 4.3. For example, constraints on the faint-end luminosity function slope, $\alpha$, could be derived from observations of dwarf galaxies with luminosities comparable to the brightest satellites we consider here; forthcoming Roman data will detect such dwarf galaxies in a range of environments. Meanwhile, the galaxy–halo luminosity scatter, $\sigma_{M}$, can be probed more directly using dwarf galaxies’ inferred dynamical masses (e.g., from stellar velocity dispersion measurements; Simon et al. 2019). Finally, the efficiency of subhalo disruption due to baryons, $\mathcal{B}$, can be better constrained using large samples of bright satellite populations (e.g., from SAGA; Geha et al. 2017; Mao et al. 2021), and priors can be tightened with improved predictions for subhalo disruption in embedded-disk simulations (Y. Wang et al., in preparation).
@@ 5 Warm Dark Matter Forecasts
Scenario A provides an optimistic forecast for the WDM constraining power of future dwarf galaxy surveys since it does not include an astrophysical signal that mimics the WDM cutoff. Scenarios B and C allow us to examine degeneracies between WDM and astrophysical cutoff parameters. For Scenario C, the WDM model we choose is disfavored by the MW satellite population as well as by other small-scale structure probes like strong gravitational lensing (Hsueh et al., 2020; Gilman et al., 2020a) and the Lyman-$\alpha$ forest (Iršič et al., 2017, 2023), and combinations thereof (e.g., Enzi et al. 2021; Nadler et al. 2021b). Nevertheless, we chose this input model so that our predicted satellite populations are significantly affected by WDM, allowing us to study degeneracies between WDM and the galaxy–halo connection in detail.

### v2

@@ 4 Galaxy Formation Forecasts
In both cases, we set $\sigma_{M}=0.2~{}\mathrm{dex}$, and $\mathcal{S}_{\mathrm{gal}}=0.2$. This value of $\sigma_{M}$ is representative of the galaxy–halo connection scatter inferred at higher masses (Wechsler & Tinker, 2018) and is near the upper limit derived from current MW satellite observations (Nadler et al., 2020b); thus, it is a helpful benchmark for our study. Meanwhile, our fiducial value of $\mathcal{S}_{\mathrm{gal}}$ yields a galaxy occupation fraction that agrees with the predictions of state-of-the-art galaxy formation models.
In particular, the “weak” cutoff scenario is broadly consistent with predictions from high-resolution simulations of galaxy formation at high redshifts (e.g., Côté et al. 2018) and from various SAMs (e.g., Kravtsov & Manwadkar 2022; Ahvazi et al. 2024). We note that the “weak” cutoff closely matches the mean galaxy occupation fraction predicted when both atomic and molecular hydrogen cooling are included in the Ahvazi et al. (2024) SAM. Meanwhile, the “strong” cutoff scenario is more consistent with predictions from hydrodynamic simulations run to lower redshifts (e.g., Sawala et al. 2016; Fitts et al. 2017; Munshi et al. 2021) and other SAMs (e.g., Benitez-Llambay & Frenk 2020). Furthermore, the “strong” cutoff is similar to (but slightly weaker than) the mean galaxy occupation fraction predicted when only atomic hydrogen cooling is included in the Ahvazi et al. (2024) SAM. We note that galaxy occupation in hydrodynamic simulations is resolution- and definition-dependent (e.g., Nadler et al. 2020b; Ahvazi et al. 2024), and that higher-resolution simulations (e.g., Wheeler et al. 2019) and convergence studies (e.g., Munshi et al. 2021) indicate that progressively lower-mass halos host “galaxies” (i.e., simulated systems with more than a critical number of star particles) as resolution increases.
@@ 4.1 Cutoff Mass Scale Posteriors
Figure 3 illustrates our results for the “strong” cutoff scenario: the left panel shows the marginalized posterior for $\mathcal{M}_{50}$ given observations of all satellites around one (blue) and two (red) MW-mass hosts, and the right panel shows the corresponding galaxy occupation fraction posteriors. We find $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})=7.9^{+0.4}_{-1.6}$ at $68\%$ confidence for one host. Thus, $\mathcal{M}_{50}$ can be probed at the $\approx 1\sigma$ level using a single complete satellite population; however, a long tail toward low $\mathcal{M}_{50}$ prevents a measurement at the $2\sigma$ level. Note that our inferred distribution of $\mathcal{M}_{50}$ is consistent with the input value of $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})=8.0$, indicating that our galaxy–halo connection reconstruction is unbiased and that our inference framework is self-consistent.
Figure 3: Left: marginalized posteriors for the galaxy formation cutoff mass scale $\mathcal{M}_{50}$ (i.e., the peak virial mass at which $50\%$ of halos host galaxies with $M_{V}<0~{}\mathrm{mag}$), from our forecasted measurements of one (blue) and two (red) complete satellite populations of MW-mass hosts. Dark (light) shaded regions show $68\%$ ($95\%$) confidence intervals. Right: marginalized posteriors for the galaxy occupation fraction, defined as the fraction of halos that host galaxies with $M_{V}<0~{}\mathrm{mag}$ as a function of peak virial mass. Dark (light) bands show the $68\%$ ($95\%$) confidence intervals from our one- and two-host forecasts. In both panels, black dashed lines show the input “strong” galaxy formation cutoff.
The tail toward low $\mathcal{M}_{50}$ arises because the “strong” cutoff primarily suppresses satellite abundances at the lowest luminosities we consider, and this effect is not significant enough to be distinguished from models that do not suppress the abundance of satellites with $M_{V}<0~{}\mathrm{mag}$ and $\mu_{V}<32~{}\mathrm{mag\ arcsec}^{-2}$. As a result, models with $\mathcal{M}_{50}\lesssim 10^{7}~{}M_{\mathrm{\odot}}$ are not distinguishable from each other and have roughly equal probability in the marginalized posterior. Our analysis of the full one-host posterior in Appendix C.1 shows that the low-$\mathcal{M}_{50}$ tail is enhanced by a degeneracy with the luminosity scatter $\sigma_{M}$. For large $\sigma_{M}$, galaxies hosted by abundant, low-mass halos preferentially up-scatter in luminosity, resulting in higher predicted satellite abundances at all luminosities. In this regime, galaxy–halo connection parameters adjust so that lower-mass halos explain more of the mock data; in turn, large values of $\mathcal{M}_{50}$ are disfavored because they prevent these low-mass halos from forming galaxies. In other words, higher-luminosity scatter increases the sensitivity of a fixed-depth survey to lower-mass subhalos, boosting the constraining power on large values of $\mathcal{M}_{50}$.
@@ 4.2 Cutoff Shape Posteriors
Figure 4: Marginalized posterior for $\mathcal{M}_{50}$ vs. $\sigma_{M}$ from our forecast using one (blue) and two (red) complete satellite populations of MW-mass hosts in the presence of a “strong” galaxy formation cutoff. Dashed lines mark the input values of $\mathcal{M}_{50}$ and $\sigma_{M}$; two-dimensional contours represent $68\%$ and $95\%$ confidence intervals. The top and side panels show marginal posteriors, where shaded regions represent $68\%$ confidence intervals. The two hosts shown by the red contour have significantly different bright satellite abundances, which markedly improves the recovery of $\sigma_{M}$ and indirectly improves $\mathcal{M}_{50}$ constraints.
@@ 4.3 Projections for Additional Hosts
As discussed in Section 3.2, it is computationally challenging to run our full MCMC forecasts for more than two hosts. Instead, we derive projected constraints on $\mathcal{M}_{50}$ by calculating the likelihood as a function of $\mathcal{M}_{50}$ with all remaining galaxy–halo connection parameters fixed to their input values. We perform this test as a function of $N_{\mathrm{hosts}}$, sampling realizations of our galaxy–halo connection model and different host combinations for each cutoff scenario and choice of $N_{\mathrm{hosts}}$. This procedure optimistically neglects galaxy–halo connection degeneracies to illustrate the ideal scaling of $\mathcal{M}_{50}$ constraints for statistically limited measurements; the resulting projections can be interpreted as cosmic variance-limited constraints from satellite populations.
@@
Figure 5 shows projected $1\sigma$ measurement uncertainties on $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ as a function of $N_{\mathrm{hosts}}$ in the “strong” cutoff scenario. We derive these uncertainties from the Fisher information evaluated at the input value of $\mathcal{M}_{50}$. For large $N_{\mathrm{hosts}}$, the $1\sigma$ uncertainty on $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ scales as $N_{\mathrm{hosts}}^{-1/2}$, as expected when adding independent measurements. With $32$ complete hosts, we project $\approx 2\%$ uncertainties on $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ at the $1\sigma$ level, assuming negligible galaxy–halo connection systematics; different host combinations affect these projected constraints at the subpercent level. For reference, there are $\approx 30$ MW-mass hosts in the Local Volume with partially characterized satellite populations (e.g., Carlsten et al. 2022). Our projections do not significantly change in the “weak” cutoff scenario, although they should only be regarded as upper limits at the $1\sigma$ level in this case.
@@
In Appendix B, we show that there is a prominent tail toward low $\mathcal{M}_{50}$ in our idealized likelihoods for the “strong” and “weak” cutoff scenarios; this tail persists for large $N_{\mathrm{hosts}}$, particularly for the “weak” cutoff. Thus, $\mathcal{M}_{50}$ is difficult to measure at the $2\sigma$ level, even in the statistically limited regime, if subhalos with $M_{\mathrm{peak}}\lesssim\mathcal{M}_{50}$ are not easily detectable. As a result, we do not expect a shallower survey than our fiducial assumption to detect either the “strong” or “weak” cutoff, even if it uses more hosts. On the other hand, constraining power qualitatively improves if satellites that do not pass our fiducial detection thresholds are included in the inference. For example, using more optimistic detection thresholds of $M_{V}<+2~{}\mathrm{mag}$ (comparable to the luminosity of the recently discovered Ursa Major III/UNIONS 1 dwarf galaxy candidate; Smith et al. 2023) and $\mu_{V}<34~{}\mathrm{mag\ arcsec}^{-2}$ (fainter than any known system and comparable to “stealth galaxies” hypothesized in previous studies; Bullock et al. 2010) disfavors models with low $\mathcal{M}_{50}$ due to nondetections of extremely faint systems. Observations of satellites below our fiducial detection thresholds can therefore improve our projected galaxy formation cutoff constraints; we leave a study of these systems in the context of our modeling framework to future work.
@@
Figure 5: Projected $1\sigma$ uncertainties for statistically limited measurements of $\log(\mathcal{M}_{50}/M_{\mathrm{\odot}})$ in the “strong” cutoff scenario, holding all remaining galaxy–halo connection parameters fixed. Points show mean measurement uncertainties as a function of the number of hosts combined in the inference for $N_{\mathrm{hosts}}=1$ (blue), $2$ (red), $4$ (orange), $8$ (green), $16$ (cyan), and $32$ (gray); the gray band shows the $68\%$ range of these uncertainties due to different host combinations. The black dashed line shows the $N_{\mathrm{hosts}}^{-1/2}$ scaling expected when combining statistically limited measurements of independent hosts.
@@ 4.4 Interpretation
Our forecasts indicate that complete measurements of two satellite populations (e.g., all satellites of the MW and M31) can probe a galaxy formation cutoff with $\mathcal{M}_{50}=10^{8}~{}M_{\mathrm{\odot}}$, at the $1\sigma$ level. Moreover, for statistically limited measurements, the uncertainty of this projected $1\sigma$ constraint shrinks as additional hosts are added. We emphasize that our forecasts only rely on halos that actually host galaxies. In particular, nondetections of faint, low surface brightness galaxies—which would primarily be hosted by halos with $M_{\mathrm{peak}}\lesssim\mathcal{M}_{50}$ in the absence of a galaxy formation cutoff—play a crucial role in our inference. Thus, modeling observational incompleteness (e.g., following Drlica-Wagner et al. 2020 and Doliva-Dolinsky et al. 2022 for the MW and M31 systems, respectively), which can mimic such nondetections, will be necessary to robustly extract a galaxy formation cutoff signal from future dwarf galaxy surveys. Upcoming facilities like Rubin will make substantial progress in this regard (Mutlu-Pakdil et al., 2021); we do not specifically consider Rubin here because careful modeling of, e.g., star–galaxy separation is needed to accurately characterize its dwarf galaxy detection sensitivity (K. Tsiane et al. 2024, in preparation).
Note that mock satellite populations in our analyses range from $\approx 100$ to $200$ objects per host (e.g., Nadler et al. 2020b). Given that subhalo and satellite statistics are approximately Poissonian (e.g., Mao et al. 2015), the statistical uncertainty of our measurements is $\approx 10\%$ per host. Because faint-end galaxy–halo connection uncertainties are much larger than $10\%$, our projected measurements are limited by galaxy–halo connection systematics rather than statistical uncertainties.
Thus, using auxiliary data to constrain galaxy–halo connection parameters can reduce uncertainties in our forecasts and help realize the statistically limited measurements derived in Section 4.3. For example, constraints on the faint-end luminosity function slope, $\alpha$, could be derived from observations of dwarf galaxies with luminosities comparable to the brightest satellites we consider here; forthcoming Roman data will detect such dwarf galaxies in a range of environments. Meanwhile, the galaxy–halo luminosity scatter, $\sigma_{M}$, can be probed more directly using dwarf galaxies’ inferred dynamical masses (e.g., from stellar velocity dispersion measurements; Simon et al. 2019). Finally, the efficiency of subhalo disruption due to baryons, $\mathcal{B}$, can be better constrained using large samples of bright satellite populations (e.g., from SAGA; Geha et al. 2017; Mao et al. 2021), and priors can be tightened with improved predictions for subhalo disruption in embedded-disk simulations (Y. Wang et al. 2024, in preparation).
@@ 5 Warm Dark Matter Forecasts
Scenario A provides an optimistic forecast for the WDM constraining power of future dwarf galaxy surveys since it does not include an astrophysical signal that mimics the WDM cutoff. Scenarios B and C allow us to examine degeneracies between WDM and astrophysical cutoff parameters. For Scenario C, the WDM model we choose is disfavored by the MW satellite population as well as by other small-scale structure probes like strong gravitational lensing (Gilman et al., 2020a; Hsueh et al., 2020) and the Lyman-$\alpha$ forest (Iršič et al., 2017, 2024), and combinations thereof (e.g., Nadler et al. 2021b; Enzi et al. 2021). Nevertheless, we chose this input model so that our predicted satellite populations are significantly affected by WDM, allowing us to study degeneracies between WDM and the galaxy–halo connection in detail.